Seattle Western University has provided the following data to be used in its service department cost allocations: Service Departments Operating Departments Facility Administration Services Programs Programs Departmental costs before allocations Student credit-hours Space occupied-square feet $2,400,000 $1,600,000 $26,800,000 $5,700,000 5,000 30,000 20,000 70,000 25,000 10,000 Required: Using the direct method, allocate the costs of the service departments to the two operating departments. Allocate the Administration cost on the basis of student credit-hours and the Facility Services cost on the basis of space occupied. (Please enter allocations from a department as negative and allocations to a department as positive. The line should add across to zero. Do not round intermediate calculations.) Service Departments Operating Departments Facility Administration Services Undergraduate Graduate rams Total rams Departmental costs before allocations 2,400,000 1,600,000 26,800,000$ 5,700,000 36,500,000 Allocations Administration costs Facility services costs Total costs after allocation $ 2,400,000$ 1,600,000$ 26,800,000$ 5,700,000$ 36,500,000
Answer #1
Service Departments Operating Departments
Administration Facility services Undergraduate Programs Graduate Programs Total
Departmental costs before allocations 2400000 1600000 26800000 5700000 36500000
Allocations:
Administration costs -2400000 0 1920000 480000 0
Facility services costs -1600000 1120000 480000 0
Total costs after allocations: 0 0 29840000 6660000 36500000
Workings:
Administration costs:
Undergraduate Programs =2400000/25000*20000
Graduate Programs =2400000/25000*5000
Facility services costs:
Undergraduate Programs =1600000/100000*70000
Graduate Programs =1600000/100000*30000